OpenStudy (anonymous):

simplify sqrt 8 - 4 sqrt 50

OpenStudy (anonymous):

\[\sqrt{8}-4\sqrt{50}\]

OpenStudy (anonymous):

ur name is kinda weird

OpenStudy (misty1212):

\[8=4\times 2\implies \sqrt8=\sqrt4\sqrt2=2\sqrt2\]

OpenStudy (anonymous):

my xbox email

OpenStudy (anonymous):

brah thats still scary

OpenStudy (misty1212):

\[50=25\times 2\implies \sqrt{50}=\sqrt{25}\sqrt2=5\sqrt2\]

OpenStudy (anonymous):

@thatsneakydad What is the square root of 8?

OpenStudy (anonymous):

so it would be \[2\sqrt{2}-5\sqrt{2}\]

OpenStudy (anonymous):

@misty1212